# High-dimensional Autoregressive Modeling for Time Series with Hierarchical Structures

### v1

@@ Abstract
High-dimensional time series often exhibit hierarchical structures represented by tensors, while statistical methodologies that can effectively exploit the structural information remain limited. We propose a supervised factor modeling framework that accommodates general hierarchical structures by extracting low-dimensional features sequentially in the mode orders that respect the hierarchical structure. Our method can select a small collection of such orders to allow for impurities in the hierarchical structures, yielding interpretable loading matrices that preserve the hierarchical relationships. A practical estimation procedure is proposed, with a hyperparameter selection scheme that identifies a parsimonious set of action orders and interim ranks, thereby revealing the possibly latent hierarchical structures. Theoretically, non‑asymptotic error bounds are derived for the proposed estimators in both regression and autoregressive settings. An application to the IPIP‑NEO‑120 personality panel illustrates superior forecasting performance and clearer structural interpretation compared with existing methods based on tensor decompositions and hierarchical factor analysis.
Keywords: Autoregressive model, Hierarchical tensor data, High-dimensional time series, Non-asymptotic properties, Supervised factor modeling
@@ 1 Introduction
High-dimensional time series abound in modern scientific and econometric studies (wang2019factor; chang2024autocovariance; chen2025estimating), which often exhibit multiway (chen2022analysis; chen2022factor; wang2024high) and partially hierarchical organizations (johnson2014measuring; gao2023divide). The data are hierarchical when they contain multiple modes representing the layers of a nested structure obtained from several groups, subgroups, and items. In many applications, however, the hierarchical structure is only partial as some modes lie outside this hierarchy or even form another hierarchy. For balanced data which admits a tensor representation, methods that can adapt to such partially hierarchical tensor time series remain limited, as existing tensor-based methods or high-dimensional dynamic factor models typically assume a fully juxtaposing or hierarchical organization (montgomery2006design; stock2011dynamic; brunner2012tutorial; yuan2023two). In practice, real datasets frequently combine both: hierarchical components accompanied by cross-cutting covariates. Developing a framework that can efficiently summarize, predict, and statistically interpret such partially hierarchical structures with complex inter-mode dependencies is therefore of both theoretical and practical importance.
A representative example is the International Personality Item Pool (IPIP) NEO-120 personality inventory (johnson2014measuring; qiao2025exploratory), which contains 120 questions assessing five broad personality traits, each with 24 questions. The data are collected over time from participants of different ages and genders, resulting in a three-way tensor time series with dimensions $12\penalty 10000\ (gender\penalty 10000\ and\penalty 10000\ age\penalty 10000\ groups)\times 5\penalty 10000\ (personalities)\times 24\penalty 10000\ (questions)$. The last two dimensions reflect a hierarchical organization, while the first does not belong to this hierarchy. Given the dataset’s scale ($12\times 5\times 24=1440$ dimensions) and the limited number of time points (approximately 100 monthly observations), there is a need for an efficient and statistically interpretable model that can extract latent features from the data for prediction of future observations. The extraction process which reduces the dimensionality in a supervised manner should respect the data’s partially hierarchical structure, when both latent feature detection and trend analysis are of sociological interest.
Natural strategies to handle data of such high dimensionality include imposing sparsity or low-rankness on the model, so that a more tractable low-dimensional structure emerges. Sparsity-inducing methods (basu2015regularized; raskutti2019convex; bertsimas2020sparse) have been widely used in high-dimensional time series analysis with its well-known prediction ability (wilms2023sparse; zheng2025interpretable). However, these methods may lack structural interpretability unless the underlying sparsity pattern aligns with a meaningful structure, as is common in genetics and neuroscience (zhou2013tensor; bi2022modeling; wu2023sparse). In IPIP-NEO-120, for instance, the responses within each personality trait are highly correlated due to the experimental design, and a sparse model may fail to capture this inherent structure.
In contrast, low-rank modeling compresses the data to capture latent structure via orthogonal loading matrices, leading to a more interpretable framework of factor modeling and analysis. Supervised Factor Modeling (SFM) (huang2025supervised) is a prominent low-rank approach that uses loading matrices to extract features (or factors) from both predictors and responses, followed by a multivariate regression on these features. Specifically for hierarchical data, literature is copious with methods that extract features tailored to the hierarchical construct, such as higher-order factor models (yung1999relationship), group factor models (klami2014group), hierarchical factor models (schmid1957development; qiao2025exact). These methods either extracts factors recursively along the hierarchy, or directly from each of its levels. However, they all assign different loading matrices to different lower-level modes under the same higher-level mode, so the number of parameters exceed the dimensions of the data. These approaches are thus only feasible in the unsupervised context, where the proportion of explained variance, rather than prediction performance, is the primary concern. When otherwise applied in the supervised context, only low dimensional settings are allowed.
As traditional low-rank methods for unsupervised factor analysis of hierarchical constructs are unsuitable for high-dimensional estimation problem, SFM methods based on tensor decomposition (huang2025supervised) are preferred given their parsimony. Among these, Tucker and canonical polyadic (CP) decompositions are widely used (kolda2009tensor). Tucker decomposition (tucker1966some) is notable for its numerical stability and interpretable orthogonal loadings, which align well with supervised factor modeling (wang2024high). However, its model complexity grows exponentially with tensor order, limiting its efficiency in high-order settings (oseledets2011tensor). This issue is exacerbated in tensor autoregressive models, where coefficient tensors often have high orders. In the IPIP-NEO-120 example, a lag-one tensor autoregressive model would involve a sixth-order coefficient tensor, and estimation based on Tucker decomposition would be computationally less efficient. Additionally, Tucker loadings are confined on Kronecker product spaces, restricting its applicability to data with complex or hierarchical structures, as discussed in Sections 2.1 and 2.2. CP decomposition (harshman1970foundations) offers a more efficient and flexible alternative (lock2018tensor), with model complexity scaling linearly with tensor order and dimensions. However, CP loadings are complicated, constraining its interpretation in unsupervised settings as a PCA surrogate (han2024cp; han2022tensor), and CP-based SFM (cai2025efficient) lacks clear interpretation for structured data.
Table 1: Comparison of different methods for supervised factor modeling. For simplicity, we assume equal dimension $p$ for each mode of the data, and equal number of features, $r$, extracted from each mode, and denote $d$ as the order of the coefficient tensor.
Models SFM based on Tensor Decomp. Factor Models for SFM
Ours CP Tucker $T$ GFM HFM HOFM
Number of parameters: $O(\cdot)$ $pr^{2}d$ prdprd $prd+r^{d}$ $pr^{2}d$ $rp^{d}$ $rp^{d}$ $rp^{d}$
Interpretation: Hierarchical $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Interpretation: Crossed $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$
Meanwhile, tensor regression based on tensor train (TT) decomposition (oseledets2011tensor) of the coefficient tensor has been proposed with its efficiency highlighted for tensors with high orders (liu2020low). Under the SFM structure, qin2025computational and si2024efficient adopt a pair of TT-styled loadings for the predictor and repsonse, and their sequential structure can capture hierarchical relationships among modes with reduced number of parameters. Specifically, each loading matrix is a product of several component matrices, with each component extracting features from one mode to which the previous mode imparts extracted features. The TT-based model is hence interpretable in the regression tasks where both the predictor and response exhibit a fully hierarchical structure (zhou2022optimal; si2024efficient). However, its strict sequential structure inherited from TT decomposition limits its applicability to data with partially hierarchical structures, as in the IPIP-NEO-120 example. Another major limitation is that the order of the component matrices in TT loadings is fixed to be forward for the predictor and backward for the response. This fixed construction hinders its application to autoregressive time series models, where the predictor and response share basically the same structure. Permuting the order of the predictor’s modes for a forced structural alignment in turn deteriorates the autoregressive model into a general stochastic regression model, complicating the temporal dependence analysis. Table 1 provides a summary of the reviewed methods under the SFM framework, comparing their number of parameters and interpretability for hierarchical and crossed (non-hierarchical) structures.
In view of the aforementioned challenges, we propose a novel tensor-based supervised factor modeling framework that effectively extracts features from partially hierarchical tensor time series, while preserving the hierarchical structure information. Our model combines admissible mode orders of the partially hierarchical structure instrinc to the data, analogous to a mixture model. Then, it extracts features sequentially, thereby capturing complex inter-mode dependencies with cross-cutting covariates obscuring any exact hierarchical structure. The proposed method is computationally efficient, as information summarization and its upward impartation to the higher-level layer bear a shared low-rank structure for the lower-level layer. This parsimony enables both accurate estimation with sample size and statistical interpretability inherited from SFM framework. Specifically, our contributions are summarized as follows:
  • We are, to the best of our knowledge, the first to bridge the methodological gap between fully hierarchical structures and non-hierarchical (juxtaposing) structures. Our proposed novel tensor-based SFM method accommodates to partially hierarchical structures in both regression and autoregression tasks, with theoretical guarantees evidenced by both simulation and empirical examples.
  • Our proposed method offers a flexible and unified way to detect from data principal and lesser hierarchical structures forming a mixture, enabled by our hyperparameter selection algorithm, when the true structure is obscured.
  • The construction in our proposed method can reveal an algebraic-graphical duality of the hierarchical structures, enshrined in an algebraic proposition, which may be of independent interest in the study of tensor decomposition and hierarchical models.
The rest of the paper is organized as follows. In Section 2, we introduce the general supervised factor modeling framework, and then present our proposed method for partially hierarchical data, followed by an interpretable tensor autoregressive model. Section 3 provides the theoretical analysis of our method and Section 4 describes its algorithms along with hyperparameter selection. In Section 5, we conduct extensive simulation studies to evaluate the performance of our method against existing alternatives. In Section 6, we apply our method to the IPIP-NEO-120 dataset for analysis. Finally, we conclude with a discussion in Section 7. All technical proofs and additional numerical results are relegated to the Supplementary Material.
Notations and preliminaries.
@@
Throughout the paper, we denote vectors by boldface small letters, e.g., $\boldsymbol{x}$; matrices by boldface capital letters, e.g., $\boldsymbol{X}$; tensors by boldface Euler capital letters, e.g., $\mathscr{X}$. For a $d$-th order tensor $\mbox{$\mathscr{X}$}\in\mathbb{R}^{p_{1}\times p_{2}\times\cdots\times p_{d}}$, denote by $\operatorname*{vec}(\mbox{$\mathscr{X}$})$, $\mbox{$\mathscr{X}$}_{(s)}$, and $\mbox{$\mathscr{X}$}]_{s}$ its vectorization, mode matricization, and sequential matricization at mode $s$ with $1\leq s\leq d$, respectively. The Frobenius norm of $\mathscr{X}$ is defined as $\left\lVert\mbox{$\mathscr{X}$}\right\rVert_{\mathrm{F}}=\sqrt{\langle\mbox{$\mathscr{X}$},\mbox{$\mathscr{X}$}\rangle}$, where $\langle\cdot,\cdot\rangle$ is the inner product. Moreover, the $(i_{1}i_{2}\ldots i_{d})$-th element of $\mathscr{X}$ is denoted by $\mbox{$\mathscr{X}$}_{i_{1}i_{2}\ldots i_{d}}$, for $i_{k}=1,2,\ldots,p_{k}$, $k=1,2,\ldots,d$. We denote $\times_{s}$ the $s$-th mode matrix product, and $\otimes$ denotes the Kronecker product. For a generic matrix $\boldsymbol{X}$, we denote by $\boldsymbol{X}^{\top}$, $\left\lVert\boldsymbol{X}\right\rVert_{\mathrm{F}}$, $\left\lVert\boldsymbol{X}\right\rVert_{2}$, and $\operatorname*{vec}(\boldsymbol{X})$ its transpose, Frobenius norm, spectral norm, and a long vector obtained by stacking all its columns, respectively. If $\boldsymbol{X}$ is further a square matrix, then denote its minimum and maximum eigenvalue by $\lambda_{\min}(\boldsymbol{X})$ and $\lambda_{\max}(\boldsymbol{X})$, respectively. Denote the orthogonal set as $\mathbb{O}^{p\times r}=\{\boldsymbol{X}\in\mathbb{R}^{p\times r}:\boldsymbol{X}^{\top}\boldsymbol{X}=\boldsymbol{I}_{r}\}$ for $p\geq r$, where $\boldsymbol{I}_{r}\in\mathbb{R}^{r\times r}$ is the identity matrix. With a little abuse of terminology, a rectangular matrix is said to be orthogonal if it belongs to $\mathbb{O}^{p\times r}$. For two real-valued sequences $x_{k}$ and $y_{k}$, $x_{k}\gtrsim y_{k}$ if there exists a $C>0$ such that $x_{k}\geq Cy_{k}$ for all $k$. In addition, we write $x_{k}\asymp y_{k}$ if $x_{k}\gtrsim y_{k}$ and $y_{k}\gtrsim x_{k}$.

### v2

# High-dimensional Autoregressive Modeling for Time Series Data with Hierarchical Structures

@@ Abstract
Modern applications have made ubiquitous high-dimensional data, especially time-dependent data, with more and more complicated structures, and it also has become more frequent to encounter the scenario of hierarchical relationships among variables. However, there is still a lack of supervised learning tool in the literature for them. To fill this gap, we introduce a new model-designing framework, and it then combines with unsupervised factor modeling tools to form an efficient and interpretable autoregressive model for high-dimensional time series with hierarchical structures. An ordinary least squares estimation is considered, and its non-asymptotic properties are established. Moreover, we propose an algorithm to search for estimates, and a boosting method is also suggested for hyperparameter selection. Simulation experiments are conducted to evaluate finite-sample performance of the proposed methodology, and its usefulness is demonstrated by an application to the Personality‑120 dataset.
Abstract
This supplementary material provides technical details and additional empirical evidence for the main paper. We first introduce the tensor notation and algebra used throughout the paper in Section A, and then present additional results on model properties in Section B, including proofs of propositions, derivations of the closed-form ALS update rules, and the SSVD procedure used for orthogonalization. We also provide complete proofs of the main theoretical guarantees in Section C, together with auxiliary lemmas in Section D. Finally, we include additional details for the simulation studies in Section E and further empirical results for the Personality-120 data analyses in Section F, which complement the findings reported in the main text.
Keywords: Autoregressive model, Hierarchical tensor data, High-dimensional time series, Non-asymptotic properties, Tensor decomposition
@@ 1 Introduction
With the rapid advancement of information technology, multi-dimensional or tensor-valued data are now ubiquitous, and particularly many of them are time-dependent. Examples can be found in medical imaging (Zhou et al., 2013; Li and Zhang, 2021), digital marketing (Bi et al., 2018; Hao et al., 2021), economics and finance (Chen et al., 2022; Wang et al., 2024) and psychology (Tang et al., 2018). It is arguably the most important task to conduct supervised learning on these data, say tensor regression (Lock, 2018; Raskutti et al., 2019) and tensor autoregression (Li and Xiao, 2021; Wang et al., 2024), while existing methods typically treat all modes in an equal manner, thus mainly for the data with fully juxtaposing structures. However, for these observed tensors, some modes may be nested within others, forming hierarchical relationships, and they can be encountered more readily for the data with a larger number of modes (Gao and Tsay, 2023; Qiao et al., 2025a). Meanwhile, it is still an open problem to model and then to predict these high-dimensional tensor-valued data, especially time-dependent data, with hierarchical structures.
As an illustrative example, we consider the Personality-120 dataset, which contains 24 questions for each of five broad personality traits, ending up with 120 questions in total. For any personality trait, each of its 24 questions is designed to quantitatively assess a specific psychological aspect, and the answers to those 24 questions jointly can be used to roughly portray an individual’s personality under this trait; see more empirical examples in psychometrics (Johnson, 2014; Qiao et al., 2025b). In addition, the dataset is collected over time from participants of different ages and genders, forming 12 demographic groups. As a result, at each time point, there is an observed three-mode tensor with dimensions of $12\penalty 10000\ (\text{demographic groups})\times 5\penalty 10000\ (\text{personality traits})\times 24\penalty 10000\ (\text{questions})$, where the last two modes reflect a hierarchical structure (questions nested within traits), while the 24 questions under each personality trait are arranged in a similar way. Moreover, the first mode (demographic groups) represents a cross-cutting covariate that does not belong to this hierarchy. It is challenging to construct an efficient time series model for the above example since it has $12\times 5\times 24=1440$ variables yet with only around 100 time points. On the other hand, to achieve an appropriate physical interpretation, another layer of difficulties is imposed due to the inherent hierarchical structure between personality traits and questions.
As arguably the most widely used model, autoregression (AR) has been a primary workhorse for many high-dimensional tasks on time series with the observed vector, matrix or tensor at each time point (Basu and Michailidis, 2015; Chen et al., 2021; Wang et al., 2024). The general AR model has $O(P^{2})$ parameters, where $P$ is the number of variables, say $P=1440$ in the Personality-120 example, and it hence is necessary to restrict the parameter space such that a reliable estimation can be achieved. A direct approach in the literature is to assume that the coefficient matrices or tensors are sparse and then to apply sparsity-inducing regularized estimation (Basu and Michailidis, 2015; Wilms et al., 2023; Zheng, 2025). The sparse AR model has been widely used, say in genetics and neuroscience (Zhou et al., 2013; Bi et al., 2022; Wu and Feng, 2023), where dependence among variables is believed to be sparse. However, for the Personality-120 example, there exists strong dependence among all $P=1440$ variables, especially each 24 variables from the same personality trait, and it is more reasonable to assume that these variables are driven by a small number of common latent factors (Qiao et al., 2025a). Similar scenario can be found for macroeconomic variables, where the factor modeling is usually considered (Lam and Yao, 2012; Bai and Wang, 2016).
Another important method for dimension reduction is to impose low-rank structures to coefficient matrices (Negahban and Wainwright, 2011) or tensors (Han et al., 2022) and, as a most commonly used tools for compressing tensors (Kolda and Bader, 2009), Tucker decomposition (Tucker, 1966) can be directly applied to the coefficient tensor of AR models (Wang et al., 2022, 2024; Huang et al., 2025). On the one hand, although the corresponding dimension of parameter spaces can be reduced dramatically, Tucker decomposition is well known not to be able to effectively compress higher-mode tensors, and it loses the efficiency very quickly as the mode of tensor increases, say greater than four or five; see Oseledets (2011). Unfortunately, consider an AR model for tensor-valued time series with mode three, and the mode of its coefficient tensor will be six. On the other hand, together with the regression setting, Tucker decomposition will lead to a nice interpretation of supervised factor modeling (Wang et al., 2024; Huang et al., 2025). Specifically, it is equivalent to first projecting responses and predictors into a small number of latent factors, termed response and predictor factors, respectively, and then there exists a regression relationship between the two types of factors. The response factors are used to summarize all predictable components, while the predictor ones contain all useful information of predictors in terms of prediction. However, the above physical interpretation will disappear for the data with hierarchical structures, such as the Personality-120 example; see Section 2.1 for details.
CP (Harshman and others, 1970) and tensor train (Oseledets, 2011) decompositions are another two popular methods for compressing tensors, and both of them are more efficient than Tucker decomposition (Kolda and Bader, 2009). However, it is difficult to interpret fitted models when CP decomposition is applied to coefficient tensors of tensor-on-tensor regression (Lock, 2018) and tensor autoregression. Tensor train (TT) decomposition may still lead to a supervised factor modeling interpretation when it is applied to tensor regression (Liu et al., 2020; Qin and Zhu, 2025) and autoregression (Si et al., 2024), while it is limited to data with a very special structure; see Remark 2 for details. To the best of our knowledge, for tensor-valued data, especially time-dependent data, with hierarchical structures, it is infeasible to achieve an efficient model with physical interpretations by directly applying an existing tensor decomposition method to coefficient tensors of the corresponding model.
Motivated by the supervised factor modeling interpretation of Tucker tensor regression, this paper introduces a new framework for constructing tensor regression and autoregressive models. Specifically, two factor modeling approaches are first applied to responses and predictors to extract features, respectively, and a low-rank structure is then imposed to the coefficient tensor so that it admits a regression relationship between response and predictor features. We may even employ different factor modeling methods for responses and predictors; see Section 2.1 for more details. Note that, for the data with general hierarchical structures, traditional factor modeling methods can be used to extract features or factors; see, for example, higher-order factor models (Yung et al., 1999), group factor models (Klami et al., 2014), and hierarchical factor models (Schmid and Leiman, 1957; Qiao et al., 2025a). These tools are unsupervised learning methods, and they admit various hierarchical interpretations by either extracting features recursively along the hierarchy, or directly from different groups of variables. We can apply them to construct tensor regression and autoregressive models within the new framework, while the loading matrices assigned to each lower-level mode are all different, leading to a huge number of parameters and hence excluding high-dimensional settings; see Section 2.2 for details. On the other hand, for many recent hierarchical data such as the Personality-120 example, they are arranged in tensor forms, i.e., each mode has the same number of variables. Moreover, the variables at each mode are also arranged in similar ways, and the corresponding factors may be formed with same loading matrices.
Table 1: Comparison of new-framework models with the higher-order factor model (HOFM), group factor model (GFM) or hierarchical factor model (HFM) being used to extract features, in terms of model complexity and interpretability for hierarchical data. The traditional framework refers to the case with tensor decomposition methods being directly applied. The coefficient tensor has $d$ modes, and each has the dimension of $p$ and rank of $r$.
Models New framework Traditional framework
Ours HOFM GFM HFM CP Tucker $T$
Number of parameters: $O(\cdot)$ $pr^{2}d$ $rp^{d}$ $rp^{d}$ $rp^{d}$ prdprd $prd+r^{d}$ $pr^{2}d$
Interpretation: Hierarchical $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$
Interpretation: Crossed $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
The first contribution of this paper is to overcome the difficulty of traditional factor modeling tools for hierarchical data by considering shared loading matrices within each layer of the hierarchy; see Table 1 for a comparison of their model complexity and interpretability for hierarchical data. Moreover, the real-world hierarchical data most likely have a mixture structure, i.e., some modes are nested, while others are crossed. We hence apply the new modeling framework to construct an efficient and interpretable regression model for these data with high dimensions in Section 2.3, and it is further extended to AR models for time-dependent data in Section 2.4. The second contribution is to consider an ordinary least squares estimation in Section 3, and its non-asymptotic properties are established. This is non-trivial since the corresponding loading matrices are non-orthogonal and, as a cost, an additional term of $\log T$ is added to the corresponding estimation error bound. Finally, Section 4 proposes an algorithm to search for the estimates, and a boosting-based method is introduced for hyperparameter selection due to the presence of hierarchical structures.
In addition, Section 5 conducts simulation experiments to evaluate numerical performance of the proposed methodology, and it is further applied to the Personality-120 dataset in Section 6. A short conclusion and discussion is given in Section 7. All technical proofs and additional numerical results are relegated to the Supplementary Material.
@@
Throughout the paper, we denote vectors by boldface small letters, e.g., $\boldsymbol{x}$; matrices by boldface capital letters, e.g., $\boldsymbol{X}$; tensors by boldface Euler capital letters, e.g., $\mathscr{X}$. For a $d$-th order tensor $\mbox{$\mathscr{X}$}\in\mathbb{R}^{p_{1}\times p_{2}\times\cdots\times p_{d}}$, denote by $\operatorname*{vec}(\mbox{$\mathscr{X}$})$, $\mbox{$\mathscr{X}$}_{(s)}$, and $\mbox{$\mathscr{X}$}]_{s}$ its vectorization, mode matricization, and sequential matricization at mode $s$ with $1\leq s\leq d$, respectively. The Frobenius norm of $\mathscr{X}$ is defined as $\left\lVert\mbox{$\mathscr{X}$}\right\rVert_{\mathrm{F}}=\sqrt{\langle\mbox{$\mathscr{X}$},\mbox{$\mathscr{X}$}\rangle}$, where $\langle\cdot,\cdot\rangle$ is the inner product. Moreover, the $(i_{1}i_{2}\ldots i_{d})$-th element of $\mathscr{X}$ is denoted by $\mbox{$\mathscr{X}$}_{i_{1}i_{2}\ldots i_{d}}$, with $1\leq k\leq d$ and $1\leq i_{k}\leq p_{k}$. We denote $\times_{s}$ the $s$-th mode matrix product, and $\otimes$ is the Kronecker product. For a generic matrix $\boldsymbol{X}$, we denote by $\boldsymbol{X}^{\top}$, $\left\lVert\boldsymbol{X}\right\rVert_{\mathrm{F}}$, $\left\lVert\boldsymbol{X}\right\rVert_{2}$, and $\operatorname*{vec}(\boldsymbol{X})$ its transpose, Frobenius norm, spectral norm, and a long vector obtained by stacking all its columns, respectively. If $\boldsymbol{X}$ is further a square matrix, we then denote its minimum and maximum eigenvalue by $\lambda_{\min}(\boldsymbol{X})$ and $\lambda_{\max}(\boldsymbol{X})$, respectively. Denote a set of orthonormal matrices by $\mathbb{O}^{p\times r}=\{\boldsymbol{X}\in\mathbb{R}^{p\times r}:\boldsymbol{X}^{\top}\boldsymbol{X}=\boldsymbol{I}_{r}\}$ for $p\geq r$, where $\boldsymbol{I}_{r}\in\mathbb{R}^{r\times r}$ is the identity matrix. For two real-valued sequences $x_{k}$ and $y_{k}$, we denote $x_{k}\gtrsim y_{k}$ or $x_{k}\lesssim y_{k}$ if there exists a $C>0$ such that $x_{k}\geq Cy_{k}$ or $x_{k}\leq Cy_{k}$ for all $k$, respectively. In addition, $x_{k}\asymp y_{k}$ if $x_{k}\gtrsim y_{k}$ and $x_{k}\lesssim y_{k}$.